OpenStudy (anonymous):

A shopkeeper gave an additional 40% discount on thereduced price after giving 25% standard concession onthatitem, if a person boughtthatitem for `1260, whatisthe original price ofthe item?

hartnn (hartnn):

let the original price be x 25% discount means, x - (25/100)x = 0.75x on this, there's 40% discount, so 0.75x - (40/100)*0.75x this amount =1260 so, you just need to solve for x in 0.75x - (40/100)*0.75x = 1260